### v1

# Hi-EF: Benchmarking Emotion Forecasting in Human-interaction

Abstract
Affective Forecasting, a research direction in psychology that predicts individuals’ future emotions, is often constrained by numerous external factors like social influence and temporal distance. To address this, we transform Affective Forecasting into a Deep Learning problem by designing an Emotion Forecasting paradigm based on two-party interactions. We propose a novel Emotion Forecasting (EF) task grounded in the theory that an individual’s emotions are easily influenced by the emotions or other information conveyed during interactions with another person. To tackle this task, we have developed a specialized dataset, Human-interaction-based Emotion Forecasting (Hi-EF), which contains 3069 two-party Multilayered-Contextual Interaction Samples (MCIS) with abundant affective-relevant labels and three modalities. Hi-EF not only demonstrates the feasibility of the EF task but also highlights its potential. Additionally, we propose a methodology that establishes a foundational and referential baseline model for the EF task and extensive experiments are provided. The dataset and code is available HERE. 33footnotetext: Engineering Research Center of Robotics, Ministry of Education, Academy for Engineering & Technology, Fudan University, Shanghai, China. 44footnotetext: School of Information Science and Technology, Fudan University, Shanghai, China.
1 Introduction
Affective computing is a branch of computer science and artificial intelligence that aims to enable computers to recognize, interpret, process, and simulate human emotions. Proposed by Rosalind Picard [1], this field combines psychology, computer science, and cognitive science to imbue machines with emotional intelligence. Current research focuses primarily on calculating and analyzing an individual’s present emotional state using facial expressions, voice, and physiological signals, which is Emotion Recognition (ER) [2, 3, 4, 5, 6, 7, 8]. However, despite significant progress in understanding and responding to current emotions, affective computing often overlooks the prediction of future emotions. Affective Forecasting [9] refers to the process by which individuals predict their future emotional reactions to various events. This cognitive mechanism plays a crucial role in decision-making, influencing choices by anticipating potential emotional outcomes. However, empirical studies [9, 10, 11] have shown that individuals often exhibit significant inaccuracies in their affective forecasts. For instance, Gilbert and Wilson [10] highlight the prevalence of “impact bias”, where people tend to overestimate the intensity and duration of their emotional responses to future events. This bias is evident in both positive and negative contexts, such as overestimating happiness from future achievements or distress from anticipated setbacks. Moreover, factors such as focalism, where individuals focus too narrowly on a single event, and underestimating emotional adaptation, contribute to these forecasting errors [11]. Understanding these predictive inaccuracies is critical, as they can lead to suboptimal decision-making in personal, professional, and social contexts. Recent advancements in this field aim to mitigate these biases by exploring the underlying cognitive and emotional mechanisms that drive affective forecasting [9]. Additionally, affective forecasting is influenced by various external factors, one of the most significant being time discounting [12, 13, 14, 15]. Time discounting refers to the tendency of individuals to devalue future events compared to immediate ones. This cognitive bias leads people to underappreciate the emotional impact of distant future events while overestimating the effect of near-term occurrences. As Loewenstein and Prelec [13] describe, individuals exhibit anomalies in intertemporal choice, leading to distorted predictions of future emotional states. Understanding the interplay between affective forecasting and time discounting is crucial for improving decision-making processes, as it helps explain why people often misjudge the long-term emotional consequences of their actions [10].
Research on affective forecasting primarily focuses on psychology and economics. However, we aim to redefine the paradigm of affective forecasting to transform this problem into a Deep Learning challenge. We propose a new task: Emotion Forecasting (EF), predicting one party’s future emotion during a two-party interaction by using short-term contextual information and the current emotional state of the other party. First, emotional exchange is bidirectional: one person’s emotions can affect the other’s, allowing for more accurate predictions of emotional changes [16]. Second, two-party interactions occur over short time frames, making the emotional predictions more reliable as responses to recent events are more direct and intense [13]. Lastly, the dynamic context of interactions provides rich background information, such as conversation content, tone, and emotional shifts, enhancing the precision of affective forecasting [17]. By remodelling affective forecasting to focus on interpersonal emotional dynamics, we can better understand and predict emotional changes in relationships, providing new tools and methods for emotion management and support [11, 10]. And we illustrate the difference between ER task and EF task in Figure 1.
Based on these, we list these criteria to construct dataset for solving EF task: (1) Multi-modal information: By analyzing facial expressions, conversational content, and vocal intonations, we can better assess a person’s emotions, thereby enabling more accurate predictions of the emotions the other party may experience. (2) Strong-Interactive Scenes: Videos should be sourced from highly interactive scenarios, enabling enough amount of interaction data for the dataset. (3) Multilayered-Contextual Information: The data format needs to include three layers of contextual information: contextual information, one party’s current emotional state, and the emotion of the other party that needs to be forecasted.
In adherence to these guidelines, we have developed the Human-interaction-based Emotion Forecasting (Hi-EF). Hi-EF boasts over 3,069 Multilayered-Contextual Interaction Samples (MCIS), featuring in excess of 5,242 video clips with abundant affective-relevant labels. To meet the stringent criteria for data collection, we have devised a comprehensive two-stage strategy to acquire a sufficient pool of candidate MCIS. Furthermore, we have implemented a reliable three-round annotation process to ensure high-quality labeling, given the complexity of the emotional information encapsulated in the labels to make sure the emotion is annotated with high quality. For the challenging EF task, we proposed the baseline method and make comprehensive experiments to demonstrate the feasibility of EF task. Besides that, more experiments of other affective computing tasks are also provided.
Figure 1: Comparing EF task with ER task.
Our work makes the following three significant contributions:
  • We have transformed Affective Forecasting from a psychological and economic problem into a Deep Learning task and developed an Emotion Forecasting paradigm based on two-party interactions. This introduces a novel task in the field of affective computing: Emotion Forecasting (EF). Previously, affective computing primarily focused on recognizing current emotions. Our work provides a new perspective for affective by aiming to predict future emotions.
  • To address this novel EF task, we have constructed a Human-interaction-based Emotion Forecasting dataset (Hi-EF). This dataset introduces a unique data format, Multilayered-Contextual Interaction Samples (MCIS), specifically designed for the EF task. MCIS encompasses short-term contextual information, one party’s current emotional state, and the other party’s future emotional state. It includes three modalities and rich affective-relevant labels for both parties. Additionally, our dataset can also be utilized for other affective computing tasks.
  • To generate high-quality MCIS, we developed a sophisticated construction procedure including candidate MCIS generation, anomaly detection and multifaceted-assisted reliable annotation.
  • We have designed a paradigm for the EF task and, based on this paradigm, established a baseline for the EF task. Extensive experiments and analyses were conducted on the Hi-EF dataset, demonstrating the feasibility of the EF task and the significance of the Hi-EF dataset. We will release our method’s code and the dataset for further research and validation.
2 Significance of EF task
Unlike traditional Affective Computing tasks that focus on recognizing current emotions, the EF task aims to predict future potential emotions based on interactional context. Accurate and reasonable predictions in this domain have a wide array of applications. We outline several future research directions and applications:
  • Individual Emotion Modeling. By analyzing extensive data from a person’s interactions, we can model their emotional responses, identifying what types of interactions are likely to elicit specific emotions. This modeling is particularly useful in the medical field for treating conditions such as depression, enabling personalized therapeutic approaches.
  • Anthropomorphic Emotion Generation. In an interactive scenario, by evaluating the emotional state and contextual information of the other party, we can predict the appropriate emotional response a human should exhibit. This ability enables the generation of anthropomorphic emotions, which is advantageous for developing emotionally intelligent robots. Such robots can respond in a more human-like and empathetic manner, thereby enhancing their effectiveness in roles that require human interaction.
3 Relevant Datasets of Emotion Forecasting: Emotion Recognition Datasets
Since the EF task is a novel task and there are no existing EF datasets, we will introduce the ER datasets that are most relevant to the EF task to aid in understanding our Hi-EF dataset. These datasets are presented in Table 1. By presenting these datasets, we aim to facilitate a better understanding of the distinctions between our Emotion Forecasting task, the Hi-EF dataset, and traditional Emotion Recognition tasks and datasets. This comparison will help clarify the unique aspects of our work. ER datasets have two mainstream types: video-driven ER datasets and multi-modal ER datasets. The primary classification is based on the dominant modalities that constitute the datasets and their corresponding related tasks, including Dynamic Facial Expression Recognition (DFER), Multi-modal Emotion Recognition (MER), and Emotion Recognition in Conversation (ERC).
Table 1: Summary of existing Affective Computing datasets (ID:1-4), ERC dataset (ID:4) and DFER datasets (ID:4-8) and our built Hi-EF. (Fac. = Facial movements; Polar. = Interaction polarity; Inten. = Interaction intensity; Emo. = Emotion; Uncer. = Uncertainty; V. = Video; A. = Audio; T. = Text.
Text Description Affective Component
ID Datasets Class Scene Fac. Scene Modality Inte. Stre. Emo. Uncer. DFER MER ERC EF
1 IEMOCAP (2008) [18] 5 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
2 HOW (2011) [19] 3 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
3 CMU-MOSEI (2018) [20] 6 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
4 MELD (2018) [21] 7 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ $\checkmark$ -
5 CAER (2019) [22] 7 1 - - V. - - $\checkmark$ - $\checkmark$ - - -
6 DFEW (2020) [23] 7 1 - - V. - - $\checkmark$ - $\checkmark$ - - -
7 FERV39K (2022) [24] 7 22 - - V. - - $\checkmark$ - $\checkmark$ - - -
8 MAFW (2022) [8] 11 1 $\checkmark$ $\checkmark$ V. &A. &T. - - $\checkmark$ - $\checkmark$ $\checkmark$ - -
9 Hi-EF (2024) 7 4 $\checkmark$ $\checkmark$ V. &A. &T. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Video-driven Emotion Recognition datasets Datasets. Video-driven Emotion Recognition datasets are dedicated to identifying human emotions from dynamic visual information. These datasets includes DFER datasets, typically originating from controlled laboratory settings, as exemplified by the CK+ [25] and Oulu-CASIA [26] datasets. Additionally, there are DFER datasets in the wild, such as AFEW [27], Aff-wild [28], AFEW-VA [29], CAER [30], DFEW [23], FERV39k [24]. Datasets like MAFW [8], which offer diverse modalities and emotion labels while still is video-driven, since the annotation is mainly based on the visual modality, while it could also be used in MER task.
Multi-modal Emotion Recognition datasets Datasets. MER datasets focuses not just the facial expression. Instead, they pay more attention to fuse the different modalities to analyze the whole emotional states of a person. And the classes is not only the expressions, but also sentiment including positive, neutral and negative and Activation-Valence-Dominance like IEMOCAP [18], HOW [19], CMU-MOSEI[20] and MELD [21]. Especially for MELD, it is text-based Emotion Recognition dataset which is designed for ERC.
The EF task differs from the ER task, although it does involve elements of ER. In the EF task, we first need to recognize one party’s emotions and then, using the contextual information, predict the potential emotions of the other party. This task structure results in the Hi-EF dataset’s MCIS format being entirely different from typical ER formats. For example, in DFER tasks, the dataset usually consists of single videos labeled with the expressions of the person in the video, typically using a seven-category classification for expressions. In contrast, MCIS includes information from both parties’ interactions. It comprises four consecutive videos: the first two videos provide contextual information, the third video presents one party’s emotional state, and the fourth video is where we need to predict the other party’s emotional state. More details about the EF task, Hi-EF, and MCIS will be elaborated in Section 4.
4 The Hi-EF Dataset
4.1 Paradigm of Emotion Forecasting
The EF task entails leveraging contextual information from the interaction between two parties and the emotions of one party to predict the potential future emotions of the other. Imagine a scenario where Party A shares a joyful experience with Party B, who shows significant interest. If A displays happiness during this interaction, it is highly probable that B will also exhibit happiness subsequently. This scenario can be divided into three layers. The first layer encompasses contextual information, serving as background and supplementary data. The second layer pertains to Party A’s emotional state, which directly influences Party B’s emotions. The third layer comprises Party B’s own emotions, which are what we aim to forecast. Therefore, we believe the EF task should incorporate the following three functions: understanding contextual information, recognizing Party A’s emotional state, and integrating these two layers of information for forecasting. These functions are defined as follows:
  • Contextual Interaction Fusion Function $I$ = $f_{CI}(\cdot)$: This function merges various layers of contextual information, providing a refined backdrop for emotion interpretation and forecast. $I$ indicates contextual information, such as non-verbal cues and the engagement levels of the participants
  • Current Party Emotion Recognition Function $E_{A}$ = $f_{ER}(\cdot)$: It assesses the current emotional state of the Party A, multiple modal information cues. $E_{A}$ is the current emotional state of Party A.
  • Future Emotion Forecasting Function $E_{B}$ = $f_{EF}(I,E_{A})$: Leveraging insights from $f_{CI}$ and $f_{ER}$, this function forecasts the probable future emotion of Party B. $E_{B}$ is the future emotion of Party B which will be forecasted.
It is important to note that in our subsequent model architecture, we do not explicitly indicate the aforementioned three functions. In our model, these functions are integrated rather than being three independent modules. Nonetheless, the model fundamentally adheres to the paradigm we designed.
Figure 2: Overview of MCIS: In order to realize the task of forecasting emotion during the interaction process, we design a new form of data MCIS with multilayered-contextual information which includes contextual information (clip I and II), current emotion state of Party A (clip III) and future emotion of Party B (clip IV). To better evaluate Party A’s emotional state and more accurately predict Party B’s emotions, we provide three modalities for MCIS: video, audio and text.
4.2 Design of Multilayered-Contextual Interaction Sample (MCIS)
To adapt to the EF task, we needed to define a data format. In the introduction, we also outlined our requirements for the dataset: Multi-modal, Strong-Interactive Scene, and Multilayered-Contextual information. Therefore, we designed the Multilayered-Contextual Interaction Sample (MCIS) as the data format for the Hi-EF dataset. Figure 2 illustrates MCIS. An MCIS primarily consists of the following components: Contextual Information: Video clip I and II; Current Information of Party A: Video clip III; Forecasted Information of Party B: Video clip IV. We provide three modalities for each video clip: video, text (utterance), and audio. Additionally, we provide annotation information for clips III and IV. It is important to note that for the Contextual Information part, we only use two clips. We did not choose more than two videos for contextual information mainly due to the following considerations: 1) In Affective Forecasting, as mentioned [12, 13, 14, 15], the time discounting issue suggests that overly long contextual information might introduce noise, affecting emotion prediction. 2) Excessively long contextual information significantly increases computational costs. 3) Contextual information serves as supplementary data, while the primary data for EF is the current emotional state of Party A [13]. This is supported by our subsequent experiments. We provide the visualization of MCIS in the Appendix.
Figure 3: Overview of the Construction Procedure for the Hi-EF Dataset. a) Shows the process of generating candidate MCIS. b) Illustrates the anomaly detection and the rules and multifaceted-assisted reliable annotation for candidate MCIS. c) Details the annotation process. All other annotation information is intended to support the more accurate labeling of emotions. Naturally, we will provide all the annotation details within the dataset.
4.3 Dataset Construction Procedure.
In this section, we present a comprehensive overview of our database creation process, which unfolds in three distinct stages: the collection of raw data and the generation of candidate video clips, the creation of candidate MCIS, and the subsequent annotation of MCIS.
Data Collection and Candidate MCIS Generation. Our dataset is sourced raw video from TV shows, rich in interactive scenes. We then extract subtitles to segment videos into clips based on the timestamp of the utterances. However, these initial clips required refinement to fit MCIS due to excess noise. We applied facial detection and a state-of-the-art ER model [31] to isolate emotionally relevant clips for MCIS construction. The purpose of this approach is to obtain video clips rich in emotions. Our methodology involves selecting three preceding clips along with the emotionally significant clip to form an MCIS unit, as detailed in Figure 3 a). And we finally get a large amount of candidate MCIS.
Anomaly Detection Rules. Our candidate MCIS generation strategy, based on utterances, occasionally results in abnormal clips. To counter this, we’ve developed a unique approach for anomaly detection to maintain annotation quality. As detailed in Figure 3 b), We defined three types of abnormalities: 1) Facial Anomalies: Clips lacking frontal facial visibility of the speaker for at least half of their duration. 2) Scene Transitions: MCIS where the final two clips are from different scenes. 3) Interaction Anomalies: MCIS featuring the same individual in the last two clips. We implemented an algorithm for initial screening against the first and third types of abnormalities, effectively identifying and discarding about half of the non-conforming MCIS. Recognizing the limitations of automated detection in our stringent criteria, manual review is essential. Therefore, we combine algorithmic and manual methods for anomaly detection to enhance accuracy, which is also contained in the next step.
Multifaceted-Assisted Reliable Annotation. The Hi-EF database stands apart due to its novel data format and undergoes a detailed three-round manual annotation process, depicted in Figure 3 (b) and (c). This process encompasses the evaluation of video, audio, and text elements. Initially, crowdsourced workers apply the abnormality criteria to sift through candidate MCIS, identifying those that are normal. These normal MCIS are then rigorously reviewed by professional annotators in the second stage. This stage is crucial for validating initial annotations and removing inconsistencies, with annotators receiving feedback to enhance their work. During this round, affective components and textual descriptions are also annotated to support reliable emotion annotation. The final round involves a thorough re-assessment of the double-checked MCIS, focusing on the last two clips of each sequence. Here, the same MCIS is annotated by five different crowdsourcing annotators. A professor, leveraging multifaceted-assisted auxiliary information, meticulously reviews these annotations to select the most appropriate or provide improved annotation. We provide extensive annotation information, including textual descriptions and emotional components. The textual descriptions cover facial actions and scene information, while the emotional components include interaction polarity, interaction intensity, emotion categories, and the degree of annotation uncertainty. These annotations are intended to comprehensively support the accuracy of emotion category labeling and can also be utilized for other tasks. Detailed annotation information will be introduced in the Appendix.
4.4 Dataset Statistics
The Hi-EF dataset has been meticulously compiled from TV shows, amassing over 100 hours of content and dividing it into more than 60,000 video clips. Following preprocessing, 30,000 clips containing emotion have been retained, and through a series of data refinement, a half of the candidates are eliminated. Manual judgment has further refined the dataset, resulting in a selection of 3,069 MCIS and 5,242 video clips with high-quality annotations. More information of dataset statistics are provided in the Appendix.
5 Experiment
5.1 Experiment Setup
[…]
Implementation Details. All experiments were implemented using PyTorch on Nvidia RTX 3090 GPUs. We set the learning rates to be in the range of 1e-2 to 1e-5, weight decay to 1e-6, and used a batch size of 8. All models were trained from scratch using optimer (Adam) for 50 epochs, with uniform frame subsampling at an interval of 8 frames. Furthermore, due to the limited availability of data in the context of Hi-EF, we employed several data augmentation techniques, including random cropping, changes in lighting conditions, and image flipping. To reduce the computational demands, all cropped images were resized to 112$\times$112, and the entire images were resized to 112$\times$168.
Evaluation Metrics. Drawing inspiration from prevalent Affective Computing tasks [32, 23, 8, 24, 33, 30, 34], we selected two traditional metrics for evaluation: Weighted Average Recall (WAR) and Unweighted Average Recall (UAR).
Figure 4: Overview of our model architecture. The first three clips in MCIS are respectively input into the intra-video fusion block, where the obtained features are then fed into the inter-video fusion block for prediction. The final prediction result is compared with the emotion label of clip IV in MCIS to calculate the loss. In the intra-video fusion block, video frames are divided into facial, posture, and scene parts, which are separately input into the vision encoder. These features are then fused to form the video modality information. Meanwhile, audio and text are input into the audio encoder and text encoder, respectively. After obtaining the features from the three modalities, they are fused to produce the feature representation of a single video. The inter-video fusion block primarily fuses the features of the three clips.
Model Architecture. Our model architecture, as depicted in Figure 4, is divided into two main parts: intra-video information fusion and inter-video information fusion. Firstly, intra-video information fusion is further divided into two segments. One segment involves the alignment and fusion of the three modalities: video, audio, and text. The other segment focuses on video modality, where we divide video frames into three parts: facial expressions, body posture, and scene information, as all three contribute to the overall emotional state of a person. For the vision encoder, we use ResNet18, I3D, and ViT. For the audio encoder, we employ the commonly used AudioCLIP Encoder, and for the text encoder, we use the CLIP text encoder. Regarding the intra-video fusion strategy, we adopt two approaches: one is the average method, and the other is fusion using a transformer. For inter-video information fusion, we primarily use average, transformer, and a combination of LSTM and transformer. It is important to note that due to the inability of ResNet18 and I3D to align with the audio encoder and text encoder, we only utilized the video modality for these two experiments.
[…]
Video Text Audio Inter Fusion WAR UAR
Face Body Scene Intra Fusion
Resnet18 Backbone
[…]
I3D Backbone
[…]
ViT Backbone
[…]
ViT ViT ViT Average CLIP AudioCLIP Average 31.44 18.55
ViT ViT ViT Transformer CLIP AudioCLIP Average 33.38 21.71
ViT ViT ViT Average CLIP AudioCLIP LSTM+Transformer 34.49 23.19
ViT ViT ViT Transformer CLIP AudioCLIP LSTM+Transformer 35.19 23.72
5.2 Experimental Results
Experiment Results of EF task. We present the experimental results of the EF task in Table 2. Our primary aim in these experiments is to evaluate the impact of different vision encoders and various intra-video and inter-video fusion strategies on the accuracy of EF. Firstly, we analyze the effect of vision encoders on the experimental results. Since ResNet18 and I3D as vision encoders cannot be aligned with the other two modalities, we only analyze the results using the video modality. We did not employ particularly complex fusion strategies for these two vision encoders, and they did not achieve superior performance. Next, we focus on the results using ViT as the vision encoder. First, we analyze the modalities. The results indicate that among the three modalities, the video modality plays the most significant role in emotion analysis. However, when the text modality is added, there is a decrease in our method’s performance. We believe this is because, when incorporating both video and text information, the text is treated as noise. However, when audio is also included, the model better understands the text information. We found that adding the text modality improves performance when there is an audio modality with video. When using all three modalities, the WAR improved by 3.82% and the UAR by 1.33% compared to using only the video modality. This indicates that we chose the right three modalities for the best performance in the EF task.
Effect of Intra-Video Fusion. During two-party interactions, emotional influence is not solely affected by facial expressions but also by body posture and the current environment. Consequently, we compared using only facial information with a combination of face, body, and scene information. The results show that the prediction accuracy using the combined face + body + scene information is higher than using facial information alone. This indicates that emotional changes during interactions are indeed influenced by these factors. Furthermore, we employed different fusion strategies to integrate these three types of information. The experimental results demonstrate that using a transformer significantly outperforms the averaging method. This is because the emotional content of the three types of information varies, requiring focused learning to capture the most relevant emotional cues.
Effect of Inter-Video Fusion. Since our MCIS includes multiple layers of information, employing appropriate fusion strategies to learn from these different layers is crucial. We used three fusion strategies: average, Rransformer, and LSTM+Transformer. The results show that the LSTM+Transformer strategy performs the best. This is because LSTM captures the temporal information between the three videos, while the transformer captures the key information that influences emotional changes across the videos. The combination of these two methods allows us to predict the other party’s potential emotions most effectively.
Effect of MCIS. We conducted ablation experiments on MCIS to compare the impact of different inputs on the EF task results which are presented in Table 3. When the input video consists of a single clip, the results show that using clip III yields the best performance. When using two video clips, the combination of the clips II and III produces the best results. This indicates that more recent information is more beneficial for the EF task. When we input three video clips, the results significantly outperform the combination of clips II and III, with a WAR improvement of 2.64% and a UAR improvement of 2.33%. This also demonstrates the effectiveness of our MCIS design.
Derived tasks of Hi-EF: In addition to the EF task, the Hi-EF dataset can also be applied to other affective computing tasks such as DFER, MER, Multilabel-classification and ERC tasks. We have included the experimental results and analyses in the Appendix.
[…]
Input Video Intra Fusion Text Audion Inter Fusion WAR UAR
I ViT Transformer CLIP AudioCLIP \ 29.50 19.38
II 30.89 20.55
III 31.44 20.80
I + II ViT Transformer CLIP AudioCLIP LSTM + Transformer 31.86 20.93
I + III 32.41 21.13
II + III 32.55 21.39
I + II + III ViT Transformer CLIP AudioCLIP LSTM + Transformer 35.19 23.72
6 Conclusions and Discussion
In this paper, we have transformed the task of Affective Forecasting from a psychological and economic problem into a Deep Learning challenge. We define a new task called Emotion Forecasting, which focuses on predicting emotions during two-party interactions. To facilitate this task, we propose the Hi-EF dataset and a baseline. The dataset features a novel data format, MCIS, which consists of multilayered-contextual interactive information between two parties. We provide rich emotion-related annotations and three modalities for each MCIS. We believe the EF task holds immense potential.
Limitations. Our baseline model does not employ a wide variety of model architectures, and the fusion strategies used are simple and commonly adopted methods without much innovation. Our primary contribution lies in designing the EF task and creating a suitable dataset and data format to accommodate this task, ultimately demonstrating its feasibility. We believe that the EF task and the Hi-EF dataset will provide a new perspective for future research in affective computing. In the future, we plan to expand the Hi-EF dataset, test more models, and continuously provide experimental results.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Abstract.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] According to our understanding, Emotion Forecasting (EF) does not appear to lead to any significant societal issues.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] The research conforms with the NeurIPS Code of Ethics.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] No theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] No theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Code is provided HERE.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We provide the implementation details in Section 5
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Statistical significance bars are not reported because it would be too computationally expensive.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We provide the implementation details in Section 5
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We cite them.
    2. (b)
      Did you mention the license of the assets? [Yes] We mentioned.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] Dataset is provided HERE .
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The videos we download and scrape from the internet are permitted to be downloaded and scraped.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] We provide the details of instructions in the supplementary materials.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [No] Crowdsourcing only annotate the video and there may be no risk.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [No] This cannot be disclosed publicly.